Theater charges 8.50 for adults and 5.50 for children. Total sales last night were 2662.50 selling 375 tickets. How many childrens tickets were sold?
Answer by mananth(16946)

adults x
child y 0

8.5 x + 5.5 y = 2662.5 .............1
1 x + 1 y = 375 .............2
Eliminate y
multiply (1)by 1
Multiply (2) by -5.5
8.5 x 5.5 y = 2662.5
-5.5 x -5.5 y = -2062.5
Add the two equations
3 x = 600
/ 3
x = 200
plug value of x in (1)
8.5 x + 5.5 y = 2662.5
1700 + 5.5 y = 2662.5
5.5 y = 2662.5 -1700
5.5 y = 962.5
y = 175
adults 200
child 175
m.ananth@hotmail.ca